OpenStudy (lisa123):

given that x=-1 is a root of g(x)=2x^3-5x^2-x+6, completely factor

OpenStudy (anonymous):

Do you know synthetic division? If you do, take the root they gave you - x = -1 - and then used that to find the other roots. Based off the 2x^3 we know that there will be 2 more roots.

OpenStudy (lisa123):

Yes I did synthetic division, and got 2x^2-7x+6 with remainder of 0

OpenStudy (anonymous):

Okay that's right, so now you should factor by either completing the square or quadratic formula

OpenStudy (lisa123):

ok well I factored by grouping and got 2x^2-7x+6 (2x^2-3x)(-4x+6) x(2x-3)-2(2x-3)

OpenStudy (lisa123):

@EmergencyExit

OpenStudy (lisa123):

@Nnesha

OpenStudy (phi):

if you think of (2x-3) as "one thing" (and it is) then x(2x-3)-2(2x-3) is like the problem xy-2y where you factor y from each term to get (x-2)y but in your problem, we factor out (2x-3) x(2x-3)-2(2x-3) = (x-2)(2x-3) that means g(x) in factored form is (x+1)(x-2)(2x-3)

OpenStudy (anonymous):

@phi You can completely simplify 2x-3 by setting it as an equation equal to 0. 2x-3=0 + 3 +3 2x = 3 x = 3/2